OpenStudy (anonymous):

Not so sure what to do.... integral of (1-sqrt(x))/(1+sqrt(x))

OpenStudy (anonymous):

yeah it does

OpenStudy (anonymous):

atm we've gone over u-sub. Not quite into trig sub nor integ. by parts.

OpenStudy (anonymous):

I'm sorry I was waiting for you to finish. But, I a bit confused as to how you balanced du and dx.

OpenStudy (dan815):

u=sqrtx du/dx=1/2 x^-(1/2) so du=1/2 x^-(1/2) dx

OpenStudy (dan815):

redoing it from here

OpenStudy (dan815):

do u follow so far?

OpenStudy (anonymous):

yes, except as to where you got that 1/2 in the final part

OpenStudy (anonymous):

\[put \sqrt{x}=u,x=u ^{2},dx=2u~du\]

OpenStudy (dan815):